Circular No. 5629

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1992 SL E. F. Helin and K. Lawrence found another fast-moving asteroidal

object on films taken Sept. 24 by Helin, Lawrence, J. Alu, and P. Rose with the Palomar 0.46-m Schmidt. Orbital elements from MPC 20935:

T = 1992 Sept. 5.320 TT  ω = 344.^{o}545
e = 0.32796  Ω = 1.086 }2000.0
q = 1.09123 AU  i = 8.476
a = 1.62377 AU n^{o} = 0.476339 P = 2.07 years
1992TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Oct. 5 1^{h}35.^{m}60 +27^{o}28.^{′}7 0.138 1.126 154.^{o}3 22.^{o}7 14.5
10 1 31.78 +31 10.2 0.153 1.138 154.6 22.1 14.8
15 1 27.97 +33 51.3 0.169 1.152 154.7 21.7 15.0
20 1 24.56 +35 41.9 0.187 1.167 154.4 21.6 15.2
25 1 21.88 +36 51.8 0.207 1.184 153.9 21.7 15.5
30 1 20.24 +37 29.7 0.229 1.201 153.0 22.0 15.8
Nov. 4 1 19.84 +37 43.8 0.252 1.220 151.8 22.6 16.0
9 1 20.70 +37 40.7 0.278 1.240 150.2 23.4 16.3
14 1 22.78 +37 25.7 0.306 1.260 148.3 24.4 16.6

SUPERNOVA 1992bb IN IRAS 21156–0747 R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of a supernova (B ∼ 17.5) in the nearly-edge-on spiral galaxy IRAS 21156– 0747. The discovery plate was taken on Sept. 30.5 UT by M. J. Drinkwater with the U.K. Schmidt Telescope. A U plate taken by Drinkwater 45 min earlier shows the supernova marginally brighter with respect to nearby stars. The supernova is not present on Sept. 2.5 to a limit of I = 19.5 on a UKST plate taken by A. Savage, but it appears at mag ∼ 18 on Sept. 27.4 on a UKST prism film taken by Drinkwater and K. M. Sealey (University of New South Wales). Drinkwater notes that the spectrum shows two emission lines, possibly [O III] at 495.9 and 500.7 nm. Coordinates measured from the B plate are α = 21^{h}15^{m}35^{s}.18, δ = −7^{o}47′42′′.5 (equinox 1950.0, uncertainty 0′′.4 in each coordinate), offset 19′′ west and 4′′ north from the galaxy’s center. No star appears in the position of the supernova on J or R survey films. A nearby star of mag 17 has end figures 32^{s}.11, 47′′.6.

1992 October 6 Daniel W. E. Green